Which of the ordered pairs is a solution for the equation 4x-2y=8?
Answer by Theo(13342)

what you need to do is x and y in the original equation to see if it holds true.

you can also solve the equation for y and then take values of x given to see if you get the y given.

start with 4x - 2y = 8
add 2y to both sides of the equation and subtract 8 from both sides of the equation to get 4x - 8 = 2y
divide both sides of the equation by 2 to get 2x - 4 = y
if x is 0, y has to be -4
if x is 1, y has to be -2
if x is 2, y has to be 0
if x is 3, y has to be 2
etc.
you can also graph the equation of y = 2x - 8

simply replacing x with the value of x and y with the value of y in the original equation to see if the equation is true might be the fastest way.

for example:
the original equation is 4x - 2y = 8
assume one of your possible solutions is (3,2)
replace x with 3 and y with 2 to get 4*3 - 2*2 = 8
simplify to get 12 - 4 = 8
simplify further to get 8 = 8
the equation is true, therefore (3,2) is a solution to the equation.